# Zhang Q Y, Xu L, Huang J H, et al

### v1

@@ 3.2.1 Uplink Distributed Receiver
When applying the cell-free wireless access network of reference [72] to LEO networks, each node is equipped with a large-scale phased array, and both satellite edge distributed units (Sat-EDU) and Sat-DU are deployed on LEO nodes. The system adopts a user-centric approach for node association, and the nodes serving users include a primary service anchor node and multiple secondary service nodes. For uplink reception, the receiver of each node independently performs multi-user detection [62]. The detected signals are sent to the primary service node Sat-DU of the user through the high-speed ISL for combined detection, then perform higher-layer physical layer processing such as demodulation and decoding. As shown in Figure 8, uplink distributed cooperation relies on ISL. Compared to the link between Sat-DU and Sat-CU, the link between Sat-EDU and Sat-DU requires the exchange of detected modulated symbol data, placing higher demands on ISL capacity.
@@ 3.2.3 Cloud-Native Distributed MIMO
In DSIN, LEO satellites move at high speed along a predetermined orbit, which continuously changes their position relative to the UTs. As a result, UTs must frequently switch links between different LEO satellites to maintain a stable network connection. This complex switching process spans two critical layers: the link and network layers. At the link layer, the primary goal is to seamlessly transfer the communication link from one satellite to another within the UTs line of sight, akin to constructing an invisible yet indispensable bridge. The network layer, however, poses a more significant challenge. During a satellite handover, when the UT connects to a new “home” satellite network, higher-level protocols such as transmission control protocol (TCP) and user datagram protocol (UDP) must migrate swiftly and seamlessly to the UTs’ new internet protocol (IP) address. This ensures that ongoing data transmissions remain uninterrupted. Unfortunately, the visible window of any single LEO satellite to a UT lasts only a few minutes, necessitating frequent handovers. This results in high signaling overhead and many issues, including reduced throughput, processing delays, data forwarding congestion, and lagging location updates. These challenges severely degrade network performance, compromise spectrum utilization, and negatively impact the user QoS.
@@
Distributed MIMO technology, which has demonstrated remarkable success in terrestrial communications, offers promising solutions for LEO satellite networks. In terrestrial-based scenarios, distributed MIMO leverages multiple access points in a collaborative, cellular-free manner to achieve high spectral efficiency, superior power efficiency, and excellent network flexibility [79]. Now, this cutting-edge technology is poised to revolutionize LEO satellite networks. By capitalizing on ultra-dense satellite constellations, ultra-fast ISLs, and line-of-sight (LoS) connectivity with terrestrial UTs, distributed MIMO is expected to unlock unprecedented communication performance. In parallel, integrating cloud-native with LEO satellite networks has injected new energy into the development of distributed MIMO systems [71]. Cloud-native platforms based virtualization infrastructure, with their robust data storage, computation, and resource integration capabilities, serve as a “cloud brain” for satellite networks [73]. When massive amounts of satellite data converge, cloud-native is able to manage a series of micro-services running on the cloud with high feasibility to achieve rapid processing and insightful analytics [80], allowing distributed MIMO systems to monitor real-time network traffic patterns and dynamically adapt to fluctuating user demands.
@@ 3.3.2 Challenges and Development of OTFS
Furthermore, researchers have recently explored the integration of MIMO-OTFS with Grant-Free Random Access (GFRA) technology for more complex multi-user detection scenarios. The authors in [121] introduced a two-dimensional pattern coupling hierarchical prior in SBL, combined with a generalized approximate MP algorithm, to improve the utilization of 2D burst block sparsity in the GFRA channel matrix, which arises from the three-dimensional structural sparsity of MIMO-OTFS in the DD domain. In [122], the investigation is extended to multi-satellite MIMO-OTFS systems, where a joint scheme for device identification, channel estimation, and symbol detection in collaborative multi-satellite GFRA is proposed, and a distributed approach is adopted to offload computational tasks to edge satellites, thereby reducing the burden on individual satellites. In addition, it is important to note that OTFS modulation ensures inherent stability of the wireless channel in the delay-Doppler (DD) domain [123, 124], which greatly facilitates the design of robust physical layer security schemes [125].
@@ 3.6 Grant-Free Random Access in Satellite-Terrestrial Networks
In NOMA scenarios, compressive sensing-based techniques for active user detection and channel estimation can effectively leverage user activity sparsity to reduce the required pilot length. To ensure optimal sparse recovery performance, designing non-orthogonal pilot sets with low correlation is crucial for minimizing user interference. A method for constructing pilot sets based on Golay sequences is proposed in [192], which successfully minimizes pilot correlation and the Peak-to-Average Power Ratio (PAPR) of spreading codes in multicarrier transmission systems. This approach provides lower PAPR while maintaining robust sparse recovery performance compared to random binary, Gaussian, pseudo-random, and Zadoff-Chu (ZC) sequences. Authors in [193] present a deterministic scheme for non-orthogonal pilot construction using a discrete Fourier transform (DFT) matrix with a mask. This method can generate $\mathcal{O}(L^{3})$ non-orthogonal sequences within a length-$L$ sequence and offers theoretical guarantees on inter-sequence correlation bounded by $\mathcal{O}(\frac{1}{\sqrt{L}})$. This enlarges the pilot set size and outperforms Gaussian random pilots by reducing collision probability. In [194], a composite preamble construction method is proposed. By combining orthogonal ZC sequences with multiple ZC root sequences with different phase rotations, this approach extends the pilot set, thereby differentiating users selecting the same orthogonal pilot.
@@
Regarding active user detection and channel estimation for LEO satellite uplink access, authors in [195] address satellite-terrestrial link channel modeling, incorporating phase shifts and channel impairments. A Bernoulli-Rayleigh message-passing algorithm is proposed, which leverages the sparse activity of ground users to reduce pilot overhead. Acknowledging the wideband transmission needs of terminal devices and the high-delay, fast-varying characteristics of satellite-terrestrial links, [61] introduces an OFDM symbol repetition mechanism. This approach addresses residual time and frequency offsets in LEO satellites, proposing an enhanced variance state propagation algorithm for joint active user detection and channel estimation. To accommodate rapidly changing satellite-terrestrial channels, [196] applies OTFS modulation for massive multi-user access, delineating a two-stage process for joint active user detection and channel estimation followed by data detection. Additionally, algorithms based on block coordinate descent, tensor Bayesian learning, and approximate message passing are also extensively employed for compressed sparse recovery problems and can be effectively utilized in uplink signal estimation and detection for TSC.
@@ 3.7.1 NOMA Enabled Multi-Satellite Multicast
Research on NOMA in satellite multicasting has seen significant progress. Early works focus on the theoretical performance of NOMA in satellite channels, demonstrating its superiority over OMA in terms of user capacity [203], spectral efficiency [204] and user fairness [202]. Subsequent studies delve into the practical aspects of implementing NOMA in satellite systems, including the design of precoding techniques to mitigate inter-beam interference [205], the development of user scheduling algorithms to optimize resource allocation [206], the improvement of information freshness to support time-critical services [207]. The application of NOMA in satellite multicasting has also been explored in the context of high-throughput satellite (HTS) systems, where the combination of NOMA with advanced beamforming techniques has been shown to significantly improve the throughput performance [208]. Moreover, the potential of NOMA in cognitive radio-inspired satellite networks has been investigated, where NOMA principles are applied to enhance spectrum sharing between primary and secondary users [209]. Recent research has also addressed the challenges of limited feedback and imperfect channel state information (CSI) in NOMA-enabled SatCom systems, proposing robust designs to ensure reliable communication in the presence of channel uncertainties [210]. Additionally, the application of NOMA in multi-beam satellite systems has been explored, aiming to improve the sum capacity throughput by using the joint beamforming and power allocation design [211]. Other studies have examined the integration of NOMA with other advanced satellite technologies, such as cognitive radio [212] and DRL [213]. Cognitive radio enables dynamic spectrum sharing between satellites and terrestrial networks, while DRL can be used to optimize resource allocation in NOMA-enabled CCS system. Over time, the focus shifted towards more complex satellite architectures, such as LEO constellations and CCS system. In particular, multi-satellite systems, where multiple satellites collaborate to provide downlink services to a common set of users, have become a key area for NOMA deployment. The transition to multi-satellite systems necessitates new approaches to address inter-satellite interference, and optimize decentralized resource allocation [214] under the constraint of asynchronous cooperative communication [215].
@@
While NOMA has shown great potential in enhancing downlink multicast communication for DSIN, several challenges remain. One of the key challenges is managing the complexity of resource allocation in DSIN. As the number of satellites in a CCS system increases, the coordination of power allocation, beam hopping, and user scheduling becomes more complex. It is necessary to develop more efficient algorithms for managing these resources in large-scale NOMA-enabled DSIN. Moreover, most of the existing PHY signal processing methods are limited to the study of two-user NOMA systems with limited satellites. Extending those solutions to large-scale networks with multiple satellites in a cluster or a swarm is of great importance. Another challenge is the management of inter-satellite interference. In CCS systems, satellites often operate in close proximity, leading to increased interference. Techniques such as multi-agent deep reinforcement learning (MADRL) enabled cooperative NOMA and beamforming can be seen as a good avenue for future research to mitigate the interference issue. In addition, as a common consumption in existing research works, perfect SIC processing may lead to overestimating the performance of NOMA-enabled CCS system, which can be addressed in future research.
@@ 3.8.2 Satellite Erasure Transmission Scheme
Further, recent studies have explored the timeliness of NC-HARQ erasure protocols, which further incorporate Age of Information (AoI) as a measure of information freshness [252]. The authors in [253] investigated the NC-HARQ with limited or no feedback retransmission for satellite communications, showing that limited retransmission optimises delay, while no retransmission yields improved AoI. An excellent erasure scheme named adaptive NC-inserted HARQ was proposed in [254] to further improve timeliness, where NC packets are dynamically inserted into the data stream based on delayed CSI to accelerate packet recovery at the receiver. The above scheme was extended to a dual-hop satellite relay scenario in [255], aiming to achieve optimal AoI and delay through a combination of adaptive CSI-based relay modes with adjustable LEC packet insertion intervals. Further enhancement on SatCom can be obtained through cross-layer optimization of error control methods. For example, the authors in [256] proposed a dual-layer coding scheme that jointly optimizes link-layer LEC and physical-layer FEC. By balancing redundancy between the LEC and FEC, i.e., finding their optimal coding rate, this erasure transfer protocol enables age-critical and reliable transmission in satellite networks. Similarly, designing a cross-layer optimization framework between the link and transport layers is another promising research direction. The benefit of this approach is that the TCP variants and LTP protocols can regard the satellite-terrestrial channel as a transparent transmission path by leveraging link-layer error control to mitigate packet loss [257]. In this case, many congestion control algorithms can be applied to the satellite network effectively, further enhancing the actual transmission performance.

### v2

@@ 3.2.1 Uplink Distributed Receiver
When applying the cell-free wireless access network of reference [83] to LEO networks, each node is equipped with a large-scale phased array, and both Sat-EDU and Sat-DU are deployed on LEO nodes. The system adopts a user-centric approach for node association, and the nodes serving users include a primary service anchor node and multiple secondary service nodes. For uplink reception, the receiver of each node independently performs multi-user detection [68]. The detected signals are sent to the primary service node Sat-DU of the user through the high-speed ISL for combined detection, then perform higher-layer physical layer processing such as demodulation and decoding. As shown in Figure 8, uplink distributed cooperation relies on ISL. Compared to the link between Sat-DU and Sat-CU, the link between Sat-EDU and Sat-DU requires the exchange of detected modulated symbol data, placing higher demands on ISL capacity.
@@ 3.2.3 Cloud-Native Distributed MIMO
In DSIN, LEO satellites move at high speed along a predetermined orbit, which continuously changes their position relative to the UTs. As a result, UTs must frequently switch links between different LEO satellites to maintain a stable network connection. This complex switching process spans two critical layers: the link and network layers. At the link layer, the primary goal is to seamlessly transfer the communication link from one satellite to another within the UTs line of sight, akin to constructing an invisible yet indispensable bridge. The network layer, however, poses a more significant challenge. During a satellite handover, when the UT connects to a new “home” satellite network, higher-level protocols such as transmission control protocol (TCP) and user datagram protocol (UDP) must migrate swiftly and seamlessly to the UTs’ new internet protocol (IP) address. This ensures that ongoing data transmissions remain uninterrupted. Unfortunately, the visible window of any single LEO satellite to a UT lasts only a few minutes, necessitating frequent handovers. This results in high signaling overhead and many issues, including reduced throughput, processing delays, data forwarding congestion, and lagging location updates. These challenges severely degrade network performance, compromise spectrum utilization, and negatively impact the user Quality of Service (QoS).
@@
Distributed MIMO technology, which has demonstrated remarkable success in terrestrial communications, offers promising solutions for LEO satellite networks. In terrestrial-based scenarios, distributed MIMO leverages multiple access points in a collaborative, cellular-free manner to achieve high spectral efficiency, superior power efficiency, and excellent network flexibility [90]. Now, this cutting-edge technology is poised to revolutionize LEO satellite networks. By capitalizing on ultra-dense satellite constellations, ultra-fast ISLs, and line-of-sight (LoS) connectivity with terrestrial UTs, distributed MIMO is expected to unlock unprecedented communication performance. In parallel, integrating cloud-native with LEO satellite networks has injected new energy into the development of distributed MIMO systems [82].
Cloud-native refers to a set of technologies that decompose applications into microservices and package them into lightweight containers for deployment and orchestration across various servers [22]. The cloud-native satellite cluster can be seen as a distributed, elastic, and horizontally scalable system, serve as a “cloud brain” for satellite networks [84]. The cloud-native satellite cluster consists of interrelated onboard microservices that isolate state in a minimal number of stateful components. As a cloud-native CCS system, it adopts practices such as microservices, containerisation, and orchestration to enable agility, scalability, and rapid development and deployment of satellite-integrated Internet applications. To set up a cloud-native distributed satellite cluster, the monolithic system must first be decomposed into self-deployable, function-specific microservices [22], which can communicate with one another through lightweight messaging protocols over ISL. Subsequently, each microservice is packaged into a container using virtualisation technologies such as Docker, KubeEdge, and Kubernetes. Each container is then orchestrated into an integrated system for functionality, with automatic configuration. Through the pooling and unified orchestration of computing power at satellite edge nodes, the cloud-native CCS system can dynamically allocate resources to each node for distributed signal processing and collaborative computation [91]. In this regard, cloud-native distributed MIMO efficiently leverages fragmented satellite node resources for cluster node status monitoring and information awareness. Based on the information gathered, it facilitates distributed collaborative transmission [92]. A cloud-based cell-free distributed massive MIMO system is investigated by [93], addressing challenges related to synchronization, calibration, and real-time baseband processing in 5G NR. General-purpose multi-core CPUs are employed for baseband signal processing.
In addition, when LEO satellite clusters perform collaborative transmission towards the ground, frequent connection handovers occur, which slow down the cluster service restart process and reduce satellite node resource recovery efficiency. On one hand, the cloud-native CCS system ensures consistency of satellite cluster information across nodes through a distributed communication protocol, while periodically updating the health status of cluster nodes to ensure the rapid dissemination and proactive exclusion of faulty node information. On the other hand, by relying on the service migration and rapid resource release and recovery capabilities of the cloud-native CCS system, the cloud-native distributed MIMO can quickly rebuild container services for new satellite nodes and initialize satellite cluster information, thereby enabling flexible and robust signal processing and satellite-terrestrial transmission.
@@ 3.3.2 Challenges and Development of OTFS
Furthermore, researchers have recently explored the integration of MIMO-OTFS with Grant-Free Random Access (GFRA) technology for more complex multi-user detection scenarios. The authors in [134] introduced a two-dimensional pattern coupling hierarchical prior in SBL, combined with a generalized approximate MP algorithm, to improve the utilization of 2D burst block sparsity in the GFRA channel matrix, which arises from the three-dimensional structural sparsity of MIMO-OTFS in the DD domain. In [135], the investigation is extended to multi-satellite MIMO-OTFS systems, where a joint scheme for device identification, channel estimation, and symbol detection in collaborative multi-satellite GFRA is proposed, and a distributed approach is adopted to offload computational tasks to edge satellites, thereby reducing the burden on individual satellites. In addition, it is important to note that OTFS modulation ensures inherent stability of the wireless channel in the DD domain [136, 137], which greatly facilitates the design of robust physical layer security schemes [138].
@@ 3.6 Grant-Free Random Access in Satellite-Terrestrial Networks
In NOMA scenarios, compressive sensing-based techniques for active user detection and channel estimation can effectively leverage user activity sparsity to reduce the required pilot length. To ensure optimal sparse recovery performance, designing non-orthogonal pilot sets with low correlation is crucial for minimizing user interference. A method for constructing pilot sets based on Golay sequences is proposed in [205], which successfully minimizes pilot correlation and the Peak-to-Average Power Ratio (PAPR) of spreading codes in multicarrier transmission systems. This approach provides lower PAPR while maintaining robust sparse recovery performance compared to random binary, Gaussian, pseudo-random, and Zadoff-Chu (ZC) sequences. Authors in [206] present a deterministic scheme for non-orthogonal pilot construction using a discrete Fourier transform (DFT) matrix with a mask. This method can generate $\mathcal{O}(L^{3})$ non-orthogonal sequences within a length-$L$ sequence and offers theoretical guarantees on inter-sequence correlation bounded by $\mathcal{O}(\frac{1}{\sqrt{L}})$. This enlarges the pilot set size and outperforms Gaussian random pilots by reducing collision probability. In [207], a composite preamble construction method is proposed. By combining orthogonal ZC sequences with multiple ZC root sequences with different phase rotations, this approach extends the pilot set, thereby differentiating users selecting the same orthogonal pilot. An iteratively constructed pilot sequence is proposed in [208] using conjugate gradient descent and space projection, subject to the constraint on the PAPR. The pilot sequence proposed in [208] can achieve up to 47% and 28% lower coherence compared with the binary Golay sequence and the traditional ZC sequence, respectively. Moreover, a Low-Correlation-Zone Periodic Sequence (LPS) is designed for the superimposed pilot structure in [209]. The average auto-correlation and cross-correlation of the LPS is $\mathcal{O}(1/L)$, which is $1/\sqrt{L}$ times lower than that of the multi-root ZC sequence, and results in a reduced decoding failure probability in satellite-based massive access systems. It is worth noting that the expansion of the non-orthogonal pilot sequence set not only alleviates access conflicts but can also be applied to pilot index modulation [210] and superimposed transmission [209], thus enhancing spectral efficiency.
@@
Regarding active user detection and channel estimation for LEO satellite uplink access, authors in [211] address satellite-terrestrial link channel modeling, incorporating phase shifts and channel impairments. A Bernoulli-Rayleigh message-passing algorithm is proposed, which leverages the sparse activity of ground users to reduce pilot overhead. Acknowledging the wideband transmission needs of terminal devices and the high-delay, fast-varying characteristics of satellite-terrestrial links, [67] introduces an OFDM symbol repetition mechanism. This approach addresses residual time and frequency offsets in LEO satellites, proposing an enhanced variance state propagation algorithm for joint active user detection and channel estimation. To accommodate rapidly changing satellite-terrestrial channels, [212] applies OTFS modulation for massive multi-user access, delineating a two-stage process for joint active user detection and channel estimation followed by data detection. However, previous studies have primarily focused on Global Navigation Satellite System (GNSS)-based solutions, which are unsuitable for power-limited Time-Sensitive Communications (TSC). In response to this, [215] proposes a joint design for device identification, channel estimation, and symbol detection in LEO satellite-enabled GFRA systems, without relying on GNSS assistance. This approach uses OTFS modulation with a message-passing-based algorithm to handle large differential delays and Doppler shifts at the satellite receiver. Additionally, algorithms based on block coordinate descent, tensor Bayesian learning, and approximate message passing are also extensively employed for compressed sparse recovery problems and can be effectively utilized in uplink signal estimation and detection for TSC.
@@
Another research branch of GRRA focuses on improving access timeliness. It is worth noting that large-scale user device access conflicts, along with access failures and the occurrence of multiple retransmissions, can lead to the accumulation of long queue buffers at IoT devices, significantly degrading the timeliness of information [216], which can be quantified using the Age of Information (AoI) metric [217]. In [218], the instantaneous AoI evolution of each device is traced through Markov analysis to derive the expression for the average AoI of the access system. A grant-free age-optimal random access protocol is then proposed to dynamically adjust the number of access slots, optimizing the system’s average AoI while enhancing the maximum throughput. Furthermore, considering that satellite networks need to accommodate service devices with diverse timeliness requirements, [219] proposes the unequal timeliness protection massive access scheme for mission critical communications in S-IoT. This scheme aims to minimize the average AoI for user groups with different timeliness demands, ensuring reliable access under their respective reliability constraints.
@@ 3.7.1 NOMA Enabled Multi-Satellite Multicast
Research on NOMA in satellite multicasting has seen significant progress. Early works focus on the theoretical performance of NOMA in satellite channels, demonstrating its superiority over orthogonal multiple access (OMA) in terms of user capacity [224], spectral efficiency [225] and user fairness [223]. Subsequent studies delve into the practical aspects of implementing NOMA in satellite systems, including the design of precoding techniques to mitigate inter-beam interference [226], the development of user scheduling algorithms to optimize resource allocation [227], the improvement of information freshness to support time-critical services [228]. The application of NOMA in satellite multicasting has also been explored in the context of high-throughput satellite (HTS) systems, where the combination of NOMA with advanced beamforming techniques has been shown to significantly improve the throughput performance [229]. Moreover, the potential of NOMA in cognitive radio-inspired satellite networks has been investigated, where NOMA principles are applied to enhance spectrum sharing between primary and secondary users [230]. Recent research has also addressed the challenges of limited feedback and imperfect CSI in NOMA-enabled SatCom systems, proposing robust designs to ensure reliable communication in the presence of channel uncertainties [231]. Additionally, the application of NOMA in multi-beam satellite systems has been explored, aiming to improve the sum capacity throughput by using the joint beamforming and power allocation design [232]. Other studies have examined the integration of NOMA with other advanced satellite technologies, such as cognitive radio [233] and DRL [234]. Cognitive radio enables dynamic spectrum sharing between satellites and terrestrial networks, while DRL can be used to optimize resource allocation in NOMA-enabled CCS system. Over time, the focus shifted towards more complex satellite architectures, such as LEO constellations and CCS system. In particular, multi-satellite systems, where multiple satellites collaborate to provide downlink services to a common set of users, have become a key area for NOMA deployment. The transition to multi-satellite systems necessitates new approaches to address inter-satellite interference, and optimize decentralized resource allocation [235] under the constraint of asynchronous cooperative communication [236].
@@
While NOMA has shown great potential in enhancing downlink multicast communication for DSIN, several challenges remain. One of the key challenges is managing the complexity of resource allocation in DSIN. As the number of satellites in a CCS system increases, the coordination of power allocation, beam hopping, and user scheduling becomes more complex. It is necessary to develop more efficient algorithms for managing these resources in large-scale NOMA-enabled DSIN. A DL-based MIMO-NOMA framework is proposed in [237] to maximize the sum data rate and energy efficiency. As an evolution of NOMA, hybrid NOMA offers superior compatibility, enabling effective coexistence between OMA and NOMA. The optimality of downlink hybrid NOMA in the two-user case has been invalidated in [238]. Furthermore, a scalable hybrid NOMA scheme is proposed in [239] to jointly control power allocation and bandwidth overlap in the presence of both imperfect CSI and SIC, thereby improving the sum rate in a spectrally efficient manner. Additionally, leveraging the spectral efficiency advantages of NOMA, Reconfigurable Intelligent Surfaces (RIS)-aided NOMA systems have garnered significant attention for maximizing the communication success ratio, while meeting the diverse QoS requirements of users and the dynamic energy constraints at the RIS [240]. In particular, the RIS-aided NOMA scheme can significantly enhance the sum rate under finite-block length coding by jointly optimising user power allocation and RIS beamforming [241]. Moreover, most of the existing PHY signal processing methods are limited to the study of two-user NOMA systems with limited satellites. Extending those solutions to large-scale networks with multiple satellites in a cluster or a swarm is of great importance. Another challenge is the management of inter-satellite interference. In CCS systems, satellites often operate in close proximity, leading to increased interference. Techniques such as multi-agent deep reinforcement learning (MADRL) enabled cooperative NOMA and beamforming can be seen as a good avenue for future research to mitigate the interference issue. In addition, as a common consumption in existing research works, perfect SIC processing may lead to overestimating the performance of NOMA-enabled CCS system, which can be addressed in future research.
@@ 3.8.2 Satellite Erasure Transmission Scheme
Further, recent studies have explored the timeliness of NC-HARQ erasure protocols, which further incorporate AoI as a measure of information freshness [217]. The authors in [278] investigated the NC-HARQ with limited or no feedback retransmission for satellite communications, showing that limited retransmission optimises delay, while no retransmission yields improved AoI. An excellent erasure scheme named adaptive NC-inserted HARQ was proposed in [279] to further improve timeliness, where NC packets are dynamically inserted into the data stream based on delayed CSI to accelerate packet recovery at the receiver. The above scheme was extended to a dual-hop satellite relay scenario in [280], aiming to achieve optimal AoI and delay through a combination of adaptive CSI-based relay modes with adjustable LEC packet insertion intervals. Further enhancement on SatCom can be obtained through cross-layer optimization of error control methods. For example, the authors in [281] proposed a dual-layer coding scheme that jointly optimizes link-layer LEC and physical-layer FEC. By balancing redundancy between the LEC and FEC, i.e., finding their optimal coding rate, this erasure transfer protocol enables age-critical and reliable transmission in satellite networks. Similarly, designing a cross-layer optimization framework between the link and transport layers is another promising research direction. The benefit of this approach is that the TCP variants and LTP protocols can regard the satellite-terrestrial channel as a transparent transmission path by leveraging link-layer error control to mitigate packet loss [282]. In this case, many congestion control algorithms can be applied to the satellite network effectively, further enhancing the actual transmission performance.